OpenStudy (anonymous):

I need to find critical points for x^5+x^2, I've taken the derivative but cannot determine the points once i've set it equal to 0

OpenStudy (anonymous):

f(x)=x^5+x^2? If so then f'(x)=5x^4+2x. If 5x^4+2x=0 then 0=x(0+/-{[0-(4)(2)]^(1/2)}/10. Maybe? Just took the quadratic formula, but it doesn't work very well...if at all.

OpenStudy (anonymous):

thanks...I confused you, I was supposed to be doing MVT anyway lol Thanks alot anyway, sorry

OpenStudy (anonymous):

Heh, no problem. =)